# The Adoption and Efficacy of Large Language Models: Evidence From Consumer Complaints in the Financial Industry

### v3

@@
Minkyu Shin Minkyu Shin: Assistant Professor of Marketing, City University of Hong Kong, minkshin@cityu.edu.hk. Jin Kim: Postdoctoral Research Associate, Northeastern University, jin.kim1@northeastern.edu. Jiwoong Shin: Professor of Marketing, Yale School of Management, jiwoong.shin@yale.edu. We thank the Department of Marketing at City University of Hong Kong for their financial support, enabling us to fully use the AI detection API. We also thank Peter Arcidiacono, Anthony Dukes, Hortense Fong, Christoph Fuchs, Soheil Ghili, Thomas L. Griffiths, Liang Guo, Vanya Klenovskiy, Yanzhi Li, Grant Packard, Stefano Puntoni, K Sudhir, Kosuke Uetake, Gal Zauberman, Joann Peck for their constructive feedback, and seminar participants at Annual Business & Generative AI Workshop by Wharton, Korea University, KAIST, INFORMS Marketing Science Conference, and Marketing Exchange Forum.    Jin Kim    Jiwoong Shin
(October 31, 2024)
Large Language Models (LLMs) are reshaping consumer decision-making, particularly in communication with firms, yet our understanding of their impact remains limited. This research explores the effect of LLMs on consumer complaints submitted to the Consumer Financial Protection Bureau from 2015 to 2024, documenting the adoption of LLMs for drafting complaints and evaluating the likelihood of obtaining relief from financial firms. Utilizing a leading AI detection tool, we analyzed over 1 million complaints and identified a significant increase in LLM usage following the release of ChatGPT. We establish a causal relationship between LLM usage and an increased likelihood of obtaining relief by employing instrumental variables to address endogeneity in LLM adoption. Experimental data further support this link, demonstrating that LLMs enhance the clarity and persuasiveness of consumer narratives. Our findings suggest that facilitating access to LLMs can help firms better understand consumer concerns and level the playing field among consumers. This underscores the importance of policies promoting technological accessibility, enabling all consumers to effectively voice their concerns.
@@ 1 Introduction
Evidence of LLM adoption among consumers varies widely according to different surveys. For example, a large-scale survey conducted between November 2023 and January 2024 indicated that ChatGPT is pervasive among various occupations in Denmark, with usage rates ranging from 79% among software developers to 34% for financial advisors Humlum and Vestergaard (2024). In contrast, a survey by the Pew Research Center in July 2023 reported a comparatively lower adoption rate of 18% among Americans, which only marginally increased to 23% by February 2024 Park and Gelles-Watnick (2023); McClain (2024). Notably, most existing studies on LLM adoption rely primarily on survey data Bick et al. (2024). Our research contributes as one of the first empirical studies to document adoption patterns using a dataset of more than 1 million consumer complaints spanning nearly a decade, from March 2015 to March 2024. Moreover, by leveraging ZIP code information in the dataset, we trace the trajectory of these adoption patterns and examine their heterogeneity across various sociodemographic groups.
Furthermore, the effectiveness of consumers using LLMs to obtain better responses from firms has yet to be empirically evaluated. While LLMs can enhance the effectiveness of communications by generating content that is coherent, informative, and formal Wu et al. (2022), they might also diminish effectiveness by producing verbose or redundant content Saito et al. (2023), or by failing to capture the subtle nuances of human emotion, which are essential for effective interpersonal communication Shahriar and Hayawi (2023). Additionally, the effective use of LLMs requires substantial domain expertise. Crafting precise prompts often demands significant experience and a thorough understanding of the specific communication context, while assessing the output also necessitates adequate knowledge of the relevant domain. For example, in the finance industry, substantial evidence of a financial illiteracy problem suggests that many consumers may lack the necessary domain knowledge required for effectively leveraging LLMs Lusardi and Mitchell (2014); Fernandes et al. (2014); Finke et al. (2017). This deficiency indicates that, even if LLMs can enhance the clarity or linguistic sophistication of consumer messages, a lack of domain knowledge may prevent consumers from using them effectively to persuade financial firms and secure favorable outcomes. Therefore, whether and to what extent LLMs impact the actual outcomes in such scenarios remains an open question.
@@
In our analysis of the Consumer Financial Protection Bureau’s extensive dataset of consumer complaints, we document empirical patterns of LLM adoption and examine how these patterns vary across sociodemographic groups. Leveraging these sociodemographic variations as instrumental variables, we provided causal evidence that LLM usage increases the likelihood of obtaining relief. To validate this causal relationship, we conducted controlled laboratory experiments, which confirm that LLM usage to refine and polish consumer complaints significantly increases the probability of receiving a favorable response. These improvements do not require consumers to possess domain-specific expertise or a deep understanding of the recipient’s preferences, as LLMs autonomously optimize key linguistic features Niven and Kao (2019); Manning et al. (2020). This process, as corroborated by our experiments, simultaneously enhances clarity, coherence, fluency, and formality Liu et al. (2019); Chen et al. (2022). This suggests that LLMs can simultaneously enhance multiple features, accommodating the heterogeneous preferences of complaint handlers and increasing the likelihood of favorable responses in line with communication accommodation theory Giles (1979). This comprehensive approach, incorporating both observational and experimental studies, demonstrates the practical benefits of LLMs in enhancing consumer advocacy through clearer and more polished messages.
@@ 2.1 Datasets & AI Detection Tool
Beyond the consumer narratives, the dataset includes a comprehensive set of structured variables that provide additional insights into the nature and resolution of the complaints. Each complaint is systematically categorized by issue type (e.g., “Improper use of your report” or “Incorrect information on your report”) and product type (e.g., “Debt collection” or “Credit reporting”). The dataset also records firm responses, categorized as “Closed with monetary relief,” “Closed with non-monetary relief,” or ”Closed with explanation,” offering a clear view of how firms address consumer grievances. In our analysis, we define a successful complaint outcome as one that results in either monetary or non-monetary relief, consistent with prior literature Dou and Roh (2019); Dou et al. (2023). [^1] This definition is particularly relevant in the financial context where distinctions between monetary and non-monetary relief are nuanced. Non-monetary relief, although not immediately quantifiable as direct financial compensation, can yield substantial economic benefits. For example, offering a foreclosure alternative like loan modification or forbearance can prevent home loss and the associated displacement costs—providing significant economic value to the consumer without direct financial payment. Detailed descriptions of all variables are provided in the Online Appendix, section 1A. Moreover, complaints are geocoded by ZIP code level, allowing for the identification of regional patterns in consumer dissatisfaction across the U.S. The ZIP code data further enables the integration of sociodemographic information from the ACS, facilitating a more granular analysis of how LLM adoption varies across different demographic and socioeconomic groups.
@@
We used the 2017–2021 ACS 5-year estimates to collect representative sociodemographic information at the ZIP code level in order to examine the heterogeneous adoption of LLMs following the release of ChatGPT in November 2022. For each ZIP code, we included key variables such as employment rate, education level, median income, total number of households, English proficiency, and internet access. These variables were chosen to capture cross-sectional variation in sociodemographic characteristics across different regions. While a detailed explanation of each variable’s calculation is provided in the Online Appendix, section 1B, we essentially followed the Census Bureau’s methodology for generating regional profiles.
@@ 2.2 Consumer Adoption of LLMs for Writing Complaints in Financial Industry
Figure 1 presents a time trend of the monthly proportion of complaints that the AI detection tool identifies as likely written by an LLM. Specifically, complaints with an AI Score $\geq$ 99% are hereafter referred to as “Likely-AI” complaints. [^3] This threshold is determined based on an examination of the AI score distribution for all complaints, which exhibits a bimodal pattern: one peak occurs above 99%, while a much more pronounced peak is observed below 5%. More details are provided in the Online Appendix, section 1D. In the months immediately following the initial release of ChatGPT (November 30, 2022), we observe a sharp increase in the proportion of Likely-AI complaints, from near 0% up to 9.8%.
@@
It is also notable that before the release of ChatGPT, the proportion of Likely-AI complaints was very close to zero. Although there were some instances where the AI detection tool incorrectly labeled a complaint as likely written by LLMs before ChatGPT was available (i.e., false positive), the rate was consistently negligible, near 0%. When applying less stringent AI Score thresholds (e.g., 80% or 70%) to define Likely-AI complaints, the rate of false positives increases. However, the surge in the share of Likely-AI complaints remains significant (see Online Appendix, section 1D). Consequently, we use a conservative threshold of 99% AI Score in our analysis to reliably identify complaints likely generated or assisted by LLMs.
@@
Note. This figure displays regional LLM adoption patterns, segmented by English proficiency quintiles derived from ACS data. Each panel corresponds to a different quintile of 5-digit ZIP codes. The top panel represents the 1st quintile, featuring regions with the lowest language barriers (4% of people with limited English proficiency on average), while the bottom panel represents the 5th quintile, with the highest (60% of people with limited English proficiency on average).
Furthermore, we observe notable regional heterogeneity in adoption patterns linked to English proficiency. For this analysis, we divided regions into five quintiles (Q1 to Q5) based on the percentage of people with limited English proficiency, using ACS data. Temporal trends were then examined separately for each quintile, as shown in Figure 1, except that the percentages are calculated based on each regional sample. As illustrated in Figure 2, each panel corresponds to one quintile, allowing us to compare the trends across regions with varying English proficiency levels. Notably, adoption was initially concentrated in regions with greater language barriers (4th and 5th quintiles), becoming more uniformly distributed across all regions by 2024, indicating that factors beyond language barriers, such as increased awareness, likely spurred this expanded adoption. These findings suggest that consumers from diverse sociodemographic backgrounds might have adopted LLMs differently when submitting complaints to financial firms. This cross-sectional variation in adoption rates served as one of the main instrumental variables in our analysis of LLM efficacy.
@@ 2.3 The Impact of Using LLMs on Getting Relief from Financial Firms
Having identified the Likely-AI complaints, we tested whether firms would respond more positively to those complaints. We categorized complaints submitted after the release of ChatGPT into two groups: Likely-AI complaints (AI scores $\geq$ 99%) and Likely-Human complaints (AI scores $\leq$ 5%). We then compared the outcome indicating whether the complaint received any form of relief — monetary or non-monetary — from firms between these groups. This comparison reveals a significant difference in the probability of receiving relief (Likely-AI: 49.3% vs. Likely-Human: 39.9%, p < .001). [^4] In comparison, the overall relief probability during the year prior to the release of ChatGPT (from December 2021 to November 2022) was 32.5%. Even after accounting for over 200 categories of complaints (where consumers select specific issues of their complaints), the positive association remains statistically significant.
While these results provide suggestive evidence of a positive association between LLM usage and the likelihood of obtaining relief, it is crucial to address potential biases that may arise from non-random selection into LLM usage. Specifically, consumers who choose to use LLMs may differ systematically from those who do not, potentially possessing unobserved characteristics that could influence firms’ responses to complaints. Such characteristics, often correlated with socioeconomic status, could bias the estimated treatment effect due to endogeneity, implying that the positive association may be less about the usage of LLMs and more about the socioeconomic status of the consumers.
@@
To address concerns regarding endogeneity and non-random selection into LLM usage, we used two demographic variables from the American Community Survey as our instrumental variables: Internet Access and English Proficiency, both measured at the ZIP code level. [^5] These two instrumental variables are also recommended by the innovative AI-assisted search method for identifying valid instrumental variables, as proposed by Han (2024). These variables are based on 5-year estimates (2017-2021), preceding the widespread release of LLMs such as ChatGPT. The timing of these data is crucial for ensuring the exogeneity of the IVs, as it mitigates concerns that they could be contemporaneously correlated with firms’ decisions to provide relief. Moreover, these IVs are likely to affect only the usage of LLM, without directly influencing relief outcomes. This exclusion restriction may hold particularly when controlling for other demographic variables, such as income. [^6] We control for a range of observable demographic characteristics such as income, education, and employment status that might otherwise confound the relationship between LLM usage and relief outcomes. By conditioning on these covariates, we aim to ensure that the IVs are not simply capturing latent socioeconomic factors that could influence both LLM adoption and firm behavior.
Our two IVs, Internet Access and English Proficiency, influence LLM adoption in opposite directions, thereby mitigating the potential impact of unobserved confounding variables. Internet Access facilitates the technical feasibility of adopting LLMs, correlating positively with higher socioeconomic factors that might independently enhance the likelihood of obtaining relief due to greater resources or knowledge. In contrast, lower English Proficiency necessitates the usage of LLMs as a tool to overcome language barriers, allowing users to compose more effective complaints. This is particularly critical in areas where socioeconomic advantages may not be as prevalent. By using these contrasting IVs, our approach effectively isolates the impact of LLM usage from potential socioeconomic confounders that might otherwise confound the observed relationships, ensuring our findings accurately reflect the direct effects of LLMs on the efficacy of consumer complaints, not just underlying socioeconomic disparities. [^7] See Mogstad et al. (2021), for a discussion on the advantages of using multiple IVs.
@@
$\vspace{-0.2cm}\ln\left(\frac{P(LLM_{i}=1)}{P(LLM_{i}=0)}\right)=\alpha_{1}Internet_{j_{i}}+\alpha_{2}English_{j_{i}}+\mathbf{X}_{j_{i}}^{\prime}\boldsymbol{\gamma}+\delta_{t_{i}}+\delta_{s_{i}}+\delta_{p_{i}}+\delta_{k_{i}}.$ (1)
For complaints submitted after the release of ChatGPT, this logistic regression models the log-odds that a given complaint $i$ was generated using an LLM, using Internet Access and English Proficiency as the primary independent variables, measured at the ZIP code level ($j_{i}$) from 2017–2021. The model includes controls for sociodemographic factors, $\mathbf{X}_{j_{i}}$, such as income, education, employment, and also accounts for several fixed effects: $\delta_{t_{i}}$ for time to capture temporal trends in LLM adoption; $\delta_{s_{i}}$ for state-level effects; $\delta_{p_{i}}$ for financial product categories (e.g., credit cards, mortgages); and $\delta_{k_{i}}$ for complaint issue types (e.g., billing, fraud). These fixed effects help isolate the exogenous influence of the IVs on LLM usage by controlling for heterogeneity across multiple dimensions.
@@
Sample Period Dec-2022 $\sim$ Mar-2024
Note. ∗∗p $<$ .01; ∗∗∗p $<$ .001
@@
Leveraging exogenous variation from two IVs in the first stage, we employ the two-stage residual inclusion (2SRI) approach outlined by Terza et al. (2008), Wooldridge (2015), and Song et al. (2020). In this 2SRI approach, which is mainly developed and employed to account for a binary outcome variable, residuals from the first-stage regression ($\hat{v}_{i}$) are used as a control function in the second-stage analysis, as shown in Equation (2). To account for the estimation uncertainty from the first-stage regression, we employ a bootstrap procedure with 1,000 replications to estimate standard errors.
$\ln\left(\frac{P(Relief_{i}=1)}{P(Relief_{i}=0)}\right)=\beta_{1}LLM_{i}+\beta_{2}\hat{v}_{i}+\mathbf{X}_{j_{i}}^{\prime}\boldsymbol{\gamma}+\delta_{t_{i}}+\delta_{s_{i}}+\delta_{j_{i}}+\delta_{p_{i}}$ (2)
The main findings, presented in Table 2, shows that consumers who used LLMs to write their complaints are more likely to receive relief from financial firms, suggesting a causal relationship. Each column in Table 2 represents the results with different residuals from various IVs used in the first stage. [^8] Our findings on the causal effect are robust to different specifications of the two-stage estimation. They remain robust when the first-stage regression is specified as a linear probability model. In addition, they are also robust to the use of generalized residuals Wooldridge (2015); Basu et al. (2018).

### v4

@@
Minkyu Shin Minkyu Shin: Assistant Professor of Marketing, City University of Hong Kong, minkshin@cityu.edu.hk. Jin Kim: Postdoctoral Research Associate, Northeastern University, jin.kim1@northeastern.edu. Jiwoong Shin: Professor of Marketing, Yale School of Management, jiwoong.shin@yale.edu.    Jin Kim    Jiwoong Shin
(February 14, 2025)
Large Language Models (LLMs) are reshaping consumer decision-making, particularly in communication with firms, yet our understanding of their impact remains limited. This research explores the effect of LLMs on consumer complaints submitted to the Consumer Financial Protection Bureau from 2015 to 2024, documenting the adoption of LLMs for drafting complaints and evaluating the likelihood of obtaining relief from financial firms. We analyzed over 1 million complaints and identified a significant increase in LLM usage following the release of ChatGPT. We find that LLM usage is associated with an increased likelihood of obtaining relief from financial firms. To investigate this relationship, we employ an instrumental variable approach to mitigate endogeneity concerns around LLM adoption. Although instrumental variables suggest a potential causal link, they cannot fully capture all unobserved heterogeneity. To further establish this causal relationship, we conducted controlled experiments, which support that LLMs can enhance the clarity and persuasiveness of consumer narratives, thereby increasing the likelihood of obtaining relief. Our findings suggest that facilitating access to LLMs can help firms better understand consumer concerns and level the playing field among consumers. This underscores the importance of policies promoting technological accessibility, enabling all consumers to effectively voice their concerns.
Competing Interests Statement: The authors declare no competing interests.
Classification: Social Science / Psychological and Cognitive Sciences
@@ 1 Introduction
Evidence of LLM adoption among consumers varies widely according to different surveys. For example, a large-scale survey conducted between November 2023 and January 2024 indicated that ChatGPT is widely utilized across diverse occupational groups in Denmark, achieving adoption rates as high as 65% among marketing specialists and journalists Humlum and Vestergaard (2025). In contrast, a survey by the Pew Research Center in July 2023 reported a comparatively lower adoption rate of 18% among Americans, which only marginally increased to 23% by February 2024 Park and Gelles-Watnick (2023); McClain (2024). Notably, most existing studies on LLM adoption rely primarily on survey data Bick et al. (2024). Our research contributes as one of the first empirical studies to document adoption patterns using a dataset of more than 1 million consumer complaints spanning nearly a decade, from March 2015 to March 2024. Moreover, by leveraging ZIP code information in the dataset, we trace the trajectory of these adoption patterns and examine their heterogeneity across various sociodemographic groups.
Furthermore, the effectiveness of consumers using LLMs to obtain better responses from firms has yet to be empirically evaluated. While LLMs can enhance the effectiveness of communications by generating content that is coherent, informative, and formal Wu et al. (2022), they might also diminish effectiveness by producing verbose or redundant content Saito et al. (2023), or by failing to capture the subtle nuances of human emotion, which are essential for effective interpersonal communication Shahriar and Hayawi (2023). Additionally, the effective use of LLMs requires substantial domain expertise. Crafting precise prompts often demands significant experience and a thorough understanding of the specific communication context, while assessing the output also necessitates adequate knowledge of the relevant domain. For example, in the finance industry, substantial evidence of a financial illiteracy problem suggests that many consumers may lack the necessary domain knowledge required for effectively leveraging LLMs Lusardi and Mitchell (2014); Fernandes et al. (2014). This deficiency indicates that, even if LLMs can enhance the clarity or linguistic sophistication of consumer messages, a lack of domain knowledge may prevent consumers from using them effectively to persuade financial firms and secure favorable outcomes. Therefore, whether and to what extent LLMs impact the actual outcomes in such scenarios remains an open question.
@@
In our analysis of the Consumer Financial Protection Bureau’s extensive dataset of consumer complaints, we document empirical patterns of LLM adoption and examine variations across sociodemographic groups. By leveraging this variation as an instrumental variable, we find evidence that LLM usage can increase the likelihood of obtaining relief. However, while instrumental variables indicate a potential causal link, they cannot fully account for all unobserved heterogeneity. To establish this causal relationship directly, we conduct controlled laboratory experiments, which confirm that LLM usage to refine and polish consumer complaints significantly increases the probability of receiving a favorable response. These improvements occur without requiring consumers to have domain-specific expertise or a deep understanding of the recipient’s preferences, as LLMs autonomously optimize key linguistic features Niven and Kao (2019); Manning et al. (2020). Our experiments demonstrate that this process simultaneously enhances clarity, coherence, fluency, and formality Liu et al. (2019); Chen et al. (2022), suggesting that LLMs can accommodate the heterogeneous preferences of complaint handlers and increase the likelihood of favorable responses in line with communication accommodation theory Giles (1979). This comprehensive approach, incorporating both observational and experimental studies, demonstrates the practical benefits of LLMs in enhancing consumer advocacy through clearer and more polished messages.
@@ 2.1 Datasets & AI Detection Tool
Beyond the consumer narratives, the dataset includes a comprehensive set of structured variables that provide additional insights into the nature and resolution of the complaints. Each complaint is systematically categorized by issue type (e.g., “Improper use of your report” or “Incorrect information on your report”) and product type (e.g., “Debt collection” or “Credit reporting”). The dataset also records firm responses, categorized as “Closed with monetary relief,” “Closed with non-monetary relief,” or “Closed with explanation,” offering a clear view of how firms address consumer grievances. In our analysis, we define a successful complaint outcome as one that results in either monetary or non-monetary relief, consistent with prior literature Dou et al. (2023). [^1] This definition is particularly relevant in the financial context where distinctions between monetary and non-monetary relief are nuanced. Non-monetary relief, although not immediately quantifiable as direct financial compensation, can yield substantial economic benefits. For example, offering a foreclosure alternative like loan modification or forbearance can prevent home loss and the associated displacement costs—providing significant economic value to the consumer without direct financial payment. Detailed descriptions of all variables are provided in the SI Appendix, section 1A. Moreover, complaints are geocoded by ZIP code level, allowing for the identification of regional patterns in consumer dissatisfaction across the U.S. The ZIP code data further enables the integration of sociodemographic information from the ACS, facilitating a more granular analysis of how LLM adoption varies across different demographic and socioeconomic groups.
@@
We used the 2017–2021 ACS 5-year estimates to collect representative sociodemographic information at the ZIP code level in order to examine the heterogeneous adoption of LLMs following the release of ChatGPT in November 2022. For each ZIP code, we included key variables such as employment rate, education level, median income, total number of households, English proficiency, and Internet access. These variables were chosen to capture cross-sectional variation in sociodemographic characteristics across different regions. While a detailed explanation of each variable’s calculation is provided in the SI Appendix, section 1B, we essentially followed the Census Bureau’s methodology for generating regional profiles.
@@ 2.2 Consumer Adoption of LLMs for Writing Complaints in Financial Industry
Figure 1 presents a time trend of the monthly proportion of complaints that the AI detection tool identifies as likely written by an LLM. Specifically, complaints with an AI Score $\geq$ 99% are hereafter referred to as “Likely-AI” complaints. [^3] This threshold is determined based on an examination of the AI score distribution for all complaints, which exhibits a bimodal pattern: one peak occurs above 99%, while a much more pronounced peak is observed below 5%. More details are provided in the SI Appendix, section 1D. In the months immediately following the initial release of ChatGPT (November 30, 2022), we observe a sharp increase in the proportion of Likely-AI complaints, from near 0% up to 9.8%.
@@
It is also notable that before the release of ChatGPT, the proportion of Likely-AI complaints was very close to zero. Although there were some instances where the AI detection tool incorrectly labeled a complaint as likely written by LLMs before ChatGPT was available (i.e., false positive), the rate was consistently negligible, near 0%.
When applying less stringent AI Score thresholds (e.g., 80% or 70%) to define Likely-AI complaints, the rate of false positives increases. However, the surge in the share of Likely-AI complaints remains significant (see SI Appendix, section 1D). Consequently, we use a conservative threshold of 99% AI Score in our analysis to reliably identify complaints likely generated or assisted by LLMs.
@@
Note. This figure displays regional variations in LLM adoption from 2020 through 2024, grouped by five English-proficiency quintiles (based on 2017–2021 ACS data) at the ZIP-code level. Each color–marker combination corresponds to a different quintile of ZIP codes (the deeper the red, the higher the proportion of individuals with limited English proficiency). For instance, the light pink line with circle markers (1st quintile) represents areas with the lowest average share of limited-English-proficient residents (around 4%), while the bold red line with diamond markers (5th quintile) indicates the highest share (around 60%). The figure demonstrates that early adoption of LLMs was more rapid in regions with a higher proportion of people with limited English proficiency.
Furthermore, we observe notable regional heterogeneity in adoption patterns linked to English proficiency. For this analysis, we divided regions into five quintiles (Q1 to Q5) based on the percentage of people with limited English proficiency, using ACS data. Temporal trends were then examined separately for each quintile, as shown in Figure 1, except that the percentages are calculated based on each regional sample. As illustrated in Figure 2, each line corresponds to one quintile, allowing us to compare the trends across regions with varying English proficiency levels. [^4] This figure focuses on the period from 2020 to 2024 to highlight adoption patterns; the complete time trend starting from 2015 is available in the SI Appendix, section 1E. Notably, adoption was initially concentrated in regions with greater language barriers (4th and 5th quintiles), becoming more uniformly distributed across all regions by 2024, indicating that factors beyond language barriers, such as increased awareness, likely spurred this expanded adoption. These findings suggest that consumers from diverse sociodemographic backgrounds might have adopted LLMs differently when submitting complaints to financial firms. This cross-sectional variation in adoption rates served as one of the main instrumental variables in our analysis of LLM efficacy.
@@ 2.3 The Impact of Using LLMs on Getting Relief from Financial Firms
Having identified the Likely-AI complaints, we tested whether firms would respond more positively to those complaints. We categorized complaints submitted after the release of ChatGPT into two groups: Likely-AI complaints (AI scores $\geq$ 99%) and Likely-Human complaints (AI scores $\leq$ 5%). We then compared the outcome indicating whether the complaint received any form of relief — monetary or non-monetary — from firms between these groups. This comparison reveals a significant difference in the probability of receiving relief (Likely-AI: 49.3% vs. Likely-Human: 39.9%, p < .001). Even after controlling for over 200 complaint issues (i.e., distinct problems indicating what each complaint is about), and time trends, this positive association remains statistically significant; the results from various regressions with different control variables are provided in SI Appendix, section 1F.
While these results provide suggestive evidence of a positive association between LLM usage and the likelihood of obtaining relief, it is crucial to address potential biases that may arise from non-random selection into LLM usage. Specifically, consumers who choose to use LLMs may differ systematically from those who do not, potentially possessing unobserved characteristics that could influence firms’ responses to complaints. Such characteristics, often correlated with socioeconomic status, could bias the estimated effect due to endogeneity, implying that the positive association may be less about the usage of LLMs and more about the socioeconomic status of the consumers.
@@
To address concerns regarding endogeneity and non-random selection into LLM usage, we used two demographic variables from the American Community Survey as our instrumental variables: Internet access and English proficiency, both measured at the ZIP code level. These variables are based on 5-year estimates (2017-2021), preceding the widespread release of LLMs such as ChatGPT. The timing of these data is crucial for ensuring the exogeneity of the IVs, as it mitigates concerns that they could be contemporaneously correlated with firms’ decisions to provide relief. Moreover, these IVs are likely to affect only the usage of LLM, without directly influencing relief outcomes. This exclusion restriction may hold particularly when controlling for other demographic variables, such as income and education. [^5] We control for a range of observable demographic characteristics such as income, education, and employment status that might otherwise confound the relationship between LLM usage and relief outcomes. By conditioning on these covariates, we aim to ensure that the IVs are not simply capturing latent socioeconomic factors that could influence both LLM adoption and firm behavior.
Specifically, our two IVs, Internet Access and English Proficiency, address different issues related to the potential endogeneity of ChatGPT usage. Internet Access captures the feasibility of adopting ChatGPT: ZIP codes with robust broadband coverage, predetermined before ChatGPT’s release, offer consumers a genuine opportunity to experiment with LLMs, whereas those lacking reliable connectivity remain effectively isolated. Because local infrastructure does not itself determine whether a complaint is granted relief, variation in Internet Access provides an exogenous source of change in LLM usage. In contrast, English Proficiency reflects the necessity of using ChatGPT: consumers with weaker English skills would face a stronger incentive to rely on AI for clarity and persuasion, a need that is largely unrelated to other drivers of success (as demonstrated in our later falsification tests). Together, these two IVs enable us to address potential confounding factors and implement a two-stage estimation that isolates the impact of LLM usage on the likelihood of obtaining relief.
@@
$\vspace{-0.2cm}\ln\left(\frac{P(LLM_{i}=1)}{P(LLM_{i}=0)}\right)=\alpha_{1}Internet_{j_{i}}+\alpha_{2}English_{j_{i}}+\mathbf{X}_{j_{i}}^{\prime}\boldsymbol{\gamma}+\delta_{t_{i}}+\delta_{s_{i}}+\delta_{p_{i}}+\delta_{k_{i}}.\vspace{0.2cm}$ (1)
For complaints submitted after the release of ChatGPT, this logistic regression models the log-odds that a given complaint $i$ was generated using an LLM (where $LLM_{i}$ is a dummy variable taking the value 1 if the complaint was generated by an LLM, and 0 otherwise); the model uses Internet access and English proficiency as the primary independent variables, measured at the ZIP code level ($j_{i}$) from 2017–2021. The model includes controls for sociodemographic factors, $\mathbf{X}_{j_{i}}$, such as income, education, employment, and a set of fixed effects to account for heterogeneity across multiple dimensions. In particular, the year-month fixed effects $\delta_{t_{i}}$ control for broad temporal trends in both LLM adoption and relief decisions, encompassing major macroeconomic shocks such as the post-COVID environment and rising inflation after 2021.666We observe an increase in consumer complaint filings beginning around January 2022—well before ChatGPT’s release. This trend underscores the importance of time-fixed effects to capture broader macroeconomic conditions that may influence complaint frequency, LLM adoption, and relief decisions. Meanwhile, the state-level fixed effects $\delta_{s_{i}}$ capture cross-state heterogeneity, the product-category fixed effects $\delta_{p_{i}}$ account for different financial products (e.g., credit cards, mortgages), and the issue-type fixed effects $\delta_{k_{i}}$ control for variation in complaint issue types (e.g., billing disputes, fraud). These fixed effects help isolate the exogenous influence of the IVs on LLM usage by controlling for heterogeneity across multiple dimensions.
@@
Sample Period Dec-2022 to Mar-2024
Note. p $<$ .05; ∗∗p $<$ .01; ∗∗∗p $<$ .001
@@
Leveraging exogenous variation from our two IVs in the first stage, we adopt the two-stage residual inclusion (2SRI) method outlined by Terza et al. (2008) and Wooldridge (2015). This approach, designed for models where both the endogenous regressor and the final outcome can be binary and estimated via a logistic link, uses the residual from the first stage as a control function. Specifically, we calculate standardized residuals ($\hat{r}_{i}\equiv\frac{LLM_{i}-\hat{P}(LLM_{i}=1)}{\sqrt{\hat{P}(LLM_{i}=1)(1-\hat{P}(LLM_{i}=1))}}$), following the idea that this can spread out the residual value to capture the rare treatment Basu et al. (2018). [^7] The main motivation for our chosen residual specification is its established suitability in 2SRI settings where the binary treatment is sparsely observed (i.e., relatively few Likely-AI complaints) Basu et al. (2018). Given the challenge of determining a priori which residual specification from a binary treatment model best captures the nonseparable error term in the outcome equation, researchers often consider alternative specifications like the raw residual, $\hat{v_{i}}\equiv LLM_{i}-\hat{P}(LLM_{i}=1)$ (equivalent to the generalized residual under a logit model) or the deviance residual Terza et al. (2008); Wooldridge (2015); Song et al. (2020). We assessed the robustness of our findings to these specifications and to the use of a linear probability model in lieu of logit, consistently confirming that LLM usage increases the likelihood of relief. Further details are provided in the SI Appendix, section 1F. We then include $\hat{r}_{i}$ in the second-stage logistic regression, as specified in Equation (2), to correct for the endogeneity of LLM usage. To account for the estimation uncertainty from the first-stage regression, we employ a bootstrap procedure with 1,000 replications to estimate standard errors.
$\ln\left(\frac{P(Relief_{i}=1)}{P(Relief_{i}=0)}\right)=\beta_{1}LLM_{i}+\beta_{2}\hat{r}_{i}+\mathbf{X}_{j_{i}}^{\prime}\boldsymbol{\gamma}+\delta_{t_{i}}+\delta_{s_{i}}+\delta_{p_{i}}+\delta_{k_{i}}\vspace{0.2cm}$ (2)
The main findings, presented in Table 2, show that consumers who used LLMs to compose their complaints are more likely to receive relief from financial firms, suggesting a potentially causal relationship. [^8] To further test the robustness of our findings, we replicate the two-stage estimation with additional company fixed effects and results are robust and we observe a very similar relationship (see SI Appendix, section 1F). Importantly, this relationship accounts for many potential confounders, including time trends, state-level cross-sectional differences, ZIP-code-level sociodemographic variables, and specific complaint issues and related products. Each column in Table 2 presents results using different residuals derived from various IVs employed in the first stage.
Since the second-stage regression employs a logit specification, the coefficients represent changes in log-odds rather than linear effects. To interpret these coefficients on the probability scale, we compute the average marginal effect of the Likely-AI dummy variable. Specifically, for each complaint, we estimate the difference in the predicted probability of obtaining relief with and without the dummy variable, holding other covariates at their observed values. We then average these differences over all observations. This procedure yields an estimated increase of approximately 10.28 percentage points in the likelihood of obtaining relief when AI is used. [^9] To further substantiate our findings, we compare complaint outcomes submitted both before and after the release of ChatGPT, focusing on regions where consumers have begun adopting LLMs (treatment group) versus those where they have not (control group). Our analysis reveals that in regions embracing LLMs, the proportion of complaints getting relief is 5 to 10 percentage points higher compared to regions where LLMs have not been widely adopted. This result supports the main conclusion of our study and is detailed in the SI Appendix, section 1F.